Terry Watkins can row about 10.6 kilometers in 1 hour downstream and 6.8 kilometers upstream in 1 hour. Find out how fast he can row in still water , and find the speed of the current.
Answer by checkley71(8403)

(10.6-6.8)/2=3.8/2=1.9 km. is the speed of the current.
10.6-1.9=8.7 km.
6.8+1.9=8.7 km. is the speed of the rowing speed.